# DT-DDNN: A Physical Layer Security Attack Detector in 5G RF Domain for CAVs

### v2

@@ V-B Experimental Validation
The experimental evaluation is conducted within the FR1 5G NR n71 band operating within the downlink frequency range of 617-652 MHz and bandwidth of 35 MHz [69]. During the initial tests, it is concluded that this spectrum is shared between Telus (with carrier frequency of ) and Rogers (with carrier frequency of ). Data acquisition is performed using ThinkRF spectrum analyzer RTSA R5500 (shown in Fig. 14) and two different types of antennas. The location of the test and therefore distance from the gNB is variable during the sampling. The experimental setup is configured with the sample rate of 15.625 MHz, the carrier frequency of , and the intermediate frequency bandwidth (IFBW) of 10 MHZ. Sampling is conducted in various environments, including indoor (behind the windows and under the desk) and outdoor (Line-of-Sight (LOS) and Non-Line-of-Sight (NLOS)) scenarios and collected samples are stored in CSV format using PyRF4 API.
To extract precise information of SSB, it is critical to perform time offset (TO) and carrier frequency offset (CFO) estimations. This is due to the lack of knowledge on the exact center frequency, which necessitates the application of a blind search. To accurately calculate the TO and CFO, we utilize the PSS correlation characteristics along with the cyclic prefix from CP-OFDM 5G waveform to fine-tune with gNB signal. The optimization problem for estimating CFO is formulated as,
@@
This work presents a robust deep-learning-based approach to detect smart and barrage jamming attacks in 5G networks with a particular focus on 5G SSB. By incorporating deep cascade learning and DWT, our DT-DDNN architecture provides remarkable accuracy in classifying a wide range of jamming scenarios, including those characterized by high SJNR values and diverse signal transmission settings. A preprocessing block is integrated to extract PSS correlation and EPNRE characteristics of the received signal which has enhanced the ability of the model to differentiate between the jammed and non-jammed observations. Inclusion of a DWT block in the model, the performance of the training process has enhanced and the training duration is reduced while maintaining the accuracy of the detection. Our results show that DT-DDNN outperforms the single threshold approach and provides more robustness and sensitivity to different jamming scenarios. DT-DDNN demonstrates improved detection probability by 10.4% and 13.2% compared to single threshold design and unprocessed IQ sample DNN design respectively. Furthermore, the adaptability of the model has been verified via several experiments with varied channel conditions, delay spread, and jamming techniques. The provided outcomes highlight the effectiveness of the suggested approach which precisely identifies jammer presence in the network with minimal false positive and miss-detection. Additionally, an experimental setup is built to assess the performance of the proposed DT-DDNN model in response to real 5G signals. The experiments conducted on the data collected from the testbed confirm the effectiveness of the system in practical scenarios. Our ongoing research includes investigation of the potential of unsupervised generative models in jamming detection to improve the performance of the system in the detection of unseen and novel jamming attacks. Furthermore, combining the CFO estimator with the design is expected to transfer the preprocessing load partially to the machine learning block.
@@ Acknowledgment
This work was supported in part by funding from the Innovation for Defence Excellence and Security (IDEaS) program from the Department of National Defence (DND). The authors would like to acknowledge the support of thinkRF in building the experimental setup for this work.

### v3

@@ V-B Experimental Validation
The experimental evaluation is conducted within the FR1 5G NR n71 band operating within the downlink frequency range of 617-652 MHz and bandwidth of 35 MHz [59]. During the initial tests, it is concluded that this spectrum is shared between Telus (with carrier frequency of ) and Rogers (with carrier frequency of ). Data acquisition is performed using ThinkRF spectrum analyzer RTSA R5500 (shown in Fig. 11), which represents CAV receiver equipment, and two different types of antennas. The location of the test and therefore distance from the gNB is variable during the sampling. The experimental setup is configured with the sample rate of 15.625 MHz, the carrier frequency of , and the intermediate frequency bandwidth (IFBW) of 10 MHZ. Sampling is conducted in various environments, including indoor (behind the windows and under the desk) and outdoor (Line-of-Sight (LOS) and Non-Line-of-Sight (NLOS)) scenarios and collected samples are stored in CSV format using PyRF4 API.
To extract precise information of SSB, it is critical to perform time offset (TO) and carrier frequency offset (CFO) estimations. This is due to the lack of knowledge of the exact center frequency, which necessitates the application of a blind search. To accurately calculate the TO and CFO, we utilize the PSS correlation characteristics along with the cyclic prefix from CP-OFDM 5G waveform to fine-tune with the gNB signal. The optimization problem for estimating CFO is formulated as,
@@
This work presents a robust deep-learning-based approach to detect smart and barrage jamming attacks in 5G networks with a particular focus on 5G SSB. By incorporating deep cascade learning and DWT, our DT-DDNN architecture provides remarkable accuracy in classifying a wide range of jamming scenarios, including those characterized by high SJNR values and diverse signal transmission settings. A preprocessing block is integrated to extract the PSS correlation and EPNRE characteristics of the received signal which has enhanced the ability of the model to differentiate between the jammed and non-jammed observations. With the inclusion of a DWT block in the model, the performance of the training process has enhanced and the training duration is reduced while maintaining the accuracy of the detection. Our results show that DT-DDNN outperforms the single threshold approach and provides more robustness and sensitivity to different jamming scenarios. DT-DDNN demonstrates improved detection probability by 10.4% and 13.2% compared to single threshold design and unprocessed IQ sample DNN design respectively. Furthermore, the adaptability of the model has been verified via several experiments with varied channel conditions, delay spread, and jamming techniques. The provided outcomes highlight the effectiveness of the suggested approach which precisely identifies jammer presence in the network with minimal false positive and miss-detection. Additionally, an experimental setup is built to assess the performance of the proposed DT-DDNN model in response to real 5G signals. The experiments conducted on the data collected from the testbed confirm the effectiveness of the system in practical scenarios. Our ongoing research includes an investigation of the potential of unsupervised generative models in jamming detection to improve the performance of the system in the detection of unseen and novel jamming attacks.
@@ Acknowledgment
This work was supported in part by funding from the Innovation for Defence Excellence and Security (IDEaS) program from the Department of National Defence (DND), and in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) CREATE TRAVERSAL Program. The authors would like to acknowledge the support of thinkRF in building the experimental setup for this work.